i need the quadratic equation to the following please
twice the square of a number is the same as eight times the same numbers

Answer by RAY100(1637)

Let n = number
.
2 n^2 = 8 n
.
2 n^2 -8 n =0
,
factor
.
2n ( n-4) = 0
.
use zero property
.
2n = 0, n=0,,,,,,,,,n-4 = 0,,, n= 4
.
check
n=0, 2* 0^2 =8*0,,,,ok
n=4, 2* 4^2 =8 * 4,,, 32=32,,,,ok